Which conic section is denoted by x = √(4-y²)
Circle,ellipse,hyperbola, parabola??

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
On squariing
 
x² = 4 - y²
 
Hence
 
x² +y² = 4
 
Hence it is circle centred at (0, 0) and has radius 2
Vikas TU
4 years ago
 
Dear student 
The standard equation of a hyperbola is x ^2/a^ 2 – y ^2/b^ 2 = 1 where b^2 = a^2(e^2 - 1).
So , iT is Hyperbola 
Hope this helps 
Good Luck